ganeshie8 (ganeshie8):

The ciphertext message produced by the RSA algorithm with \(\large (n,~k) = (1643,~223)\) is \[\large \text{0833 0823 1130 0055 0329 1099} \] Break the cipher first then decrypt the above message

OpenStudy (freckles):

lol can we cheat and have wolfram factor 1643

ganeshie8 (ganeshie8):

ofcourse yes! after all RSA is all about making things computationally infeasible..

OpenStudy (freckles):

n=p*q=31*53 totient(n)=30*52=1560 ed congrent 1 (mod 1560) e is the public key so I take your k to be my e 223d congruent 1 (mod 1560) we can solve this for the private key d

OpenStudy (freckles):

\[223d-1=1560k \\ 223-1560k=1 \\ 1560=223(6)+222 \\ 223=222(1)+1 \\ 223-222=1 \\ 223-(1560-223(6))=1 \\ 7(223)-1560=1\]

OpenStudy (freckles):

\[d=7\]

OpenStudy (freckles):

\[m=c^d (\mod n) \\ m=c^7 (\mod 1643)\]

OpenStudy (freckles):

is the decryption function

ganeshie8 (ganeshie8):

Wow! I was thinking it would be tough but you're making it look too simple haha

ganeshie8 (ganeshie8):

decryption key = 7 is right !

OpenStudy (freckles):

0833 this is one member of the ciphertext right ? \[m=833^7 (\mod 1643)\]

ganeshie8 (ganeshie8):

Yes and yes !

ganeshie8 (ganeshie8):

we may use wolfram to reduce that i guess

OpenStudy (freckles):

sorry my computer wasn't loading this thingy

OpenStudy (freckles):

yeah I don't know i non-wolfram way to compute that last thingy i said

OpenStudy (freckles):

like I would use repeated squares but 833^2 is ugly

ganeshie8 (ganeshie8):

using calculator is fine, yeah there is no fun in doing it manually

OpenStudy (freckles):

\[c=0833 \text{ } 0823 \text{ } 1130 \text{ } 0055 \text{ } 0329 \text{ } 1099 \\ m=170 \text{ } 415 \text{ } 112 \text{ } 499 \text{ } 131 \text{ } 422 \]

ganeshie8 (ganeshie8):

I got the same xD i am using below site to convert numbers to alphabets http://cryptoclub.math.uic.edu/substitutioncipher/ltr_num.htm

ganeshie8 (ganeshie8):

170 415 112 499 131 422 grouping 2 each and replacing 99 with space : ``` 17 04 15 11 24 13 14 22 ```

OpenStudy (freckles):

17=r 04=e 15=p 11=l 24=y 13=n 14=o 22=w reply now you made this up?

ganeshie8 (ganeshie8):

Nice :) Nope it is a textbook exercise.. Honestly I didn't expect it to be too easy lol ty :)

OpenStudy (freckles):

well I kinda looked at the RSA some back in college and then also a calculator helps :p

OpenStudy (freckles):

i made something like this a long time ago i will see if i can find it

ganeshie8 (ganeshie8):

the textbook encourages to use calculator for all exercises in cryptography please id like to very much go over all interestig crypto problems xD

OpenStudy (freckles):

ciphertext: 15 02 25 17 25 , 01 17 25 , 52 03 , 11 04 09 49 24, 05 41 , 36 25 05 36 23 25 , 47 , 15 02 05 24 25 , 12 02 05 , 42 01 09 , 21 09 49 25 17 24 15 01 09 49 , 18 04 09 01 17 20 , 01 09 49 , 15 02 05 24 25 , 42 01 09 , 42 01 09 32 15 , 46 now "words" are separated by , words can be just a punctuation mark m=01-26 is the normal A-Z m= 27 you have 00 m=28 you have 01 m=38 you have : m=41 yu have . m=43 you have '

OpenStudy (freckles):

I'm trying to piece what the origial problem was based on pieces

OpenStudy (freckles):

lol I still haven't gave you public key yet

OpenStudy (freckles):

(n,e)=(55,27) is what is published to the public

ganeshie8 (ganeshie8):

yes im waiting haha and that code is pretty common i think... my textbook uses the same code..

OpenStudy (freckles):

Now you are also suppose to find any type-o 's in the above sentence. Don't worry there isn't many.

OpenStudy (freckles):

But the sentence above is a computer science joke.

ganeshie8 (ganeshie8):

Let me try and find the decryption key first n=p*q=5*11 totient(n)=4*10=40 ed congrent 1 (mod 55) e is the public key so I take your k to be my e 27d congruent 1 (mod 55) we can solve this for the private key d

ganeshie8 (ganeshie8):

\(27d \equiv 1 \pmod{55} \implies d \equiv \dfrac{1}{27}\equiv \dfrac{2}{54}\equiv \dfrac{2}{-1}\equiv -2\equiv 53 \pmod{55}\) so \(d = 53\)

OpenStudy (freckles):

oh wait

OpenStudy (freckles):

ed congruent 1 (mdd 40) right?

OpenStudy (freckles):

mod*

ganeshie8 (ganeshie8):

Oops yes my mistake~ \[27d \equiv 1 \pmod{40} \implies d \equiv \dfrac{1}{27}\equiv \dfrac{1}{-13} \equiv \dfrac{-3}{39}\equiv \dfrac{-3}{-1}\equiv 3\pmod{40} \] so \(d = 3\) ?

ganeshie8 (ganeshie8):

c = (15 02 25 17 25 01 17 25 52 03 11 04 09 49 24 05 41 36 25 05 36 23 25 47 15 02 05 24 25 12 02 05 42 01 09 21 09 49 25 17 24 15 01 09 49 18 04 09 01 17 20 01 09 49 15 02 05 24 25 42 01 09 42 01 09 32 15 46) m = (20 8 5 18 5 1 18 5 28 27 11 9 14 4 19 15 6 16 5 15 16 12 5 38 20 8 15 19 5 23 8 15 3 1 14 21 14 4 5 18 19 20 1 14 4 2 9 14 1 18 25 1 14 4 20 8 15 19 5 3 1 14 3 1 14 43 20 41)

ganeshie8 (ganeshie8):

THERE ARE 10 KINDS OF PEOPLE: THOSE WHO CAN UNDERSTAND BINARY AND THOSE CAN'T. i heard this before nice one ;) http://archive.oreilly.com/pub/post/there_are_10_kinds_of_people_e.html

OpenStudy (freckles):

lol